OpenStudy (anonymous):

SOMEONE PLEASE HELP ME! will give medal and fan! Ice Cream: An ice-cream cone has a radius of 2.3 cm and a height of 6.2 cm. The cone is filled with ice-cream and one additional scoop in the shape of a sphere with the same radius as the cone is placed on top. What is the amount of ice cream in cubic cm? Round to the nearest hundredth. There are______ cubic centimeters of ice cream

OpenStudy (anonymous):

@shiraz14 could u help quick?

OpenStudy (shiraz14):

@katlin95 : It's the volume of the cone added to the volume of the sphere. Can you work out each of these volumes independently then add them?

OpenStudy (anonymous):

no, not really. Ive been working on this problem for awhile.

OpenStudy (anonymous):

can you give me the formula to use, so that I am able to find my answer?

OpenStudy (shiraz14):

Volume of cone = (1/3)*(πr^2h) Volume of sphere = (4/3)*(πr^3)

OpenStudy (anonymous):

Okay, so where would I put the 2.3 and the 6.2?

OpenStudy (shiraz14):

r=2.3; h=6.2

OpenStudy (anonymous):

I got 34.32 for the volume of the cone, is that right?

OpenStudy (shiraz14):

I got 34.34

OpenStudy (anonymous):

ughh stupid thing exited out of me and gave me a different question... this is the new question: Ice Cream: An ice-cream cone has a radius of 2.2 cm and a height of 6.2 cm. The cone is filled with ice-cream and one additional scoop in the shape of a sphere with the same radius as the cone is placed on top. What is the amount of ice cream in cubic cm? Round to the nearest hundredth. There are cubic centimeters of ice cream.